OpenStudy (anonymous):

Find the probability. A die is rolled 20 times and the number of twos that come up is tallied. Find the probability of getting exactly five twos. A. 0.129 B. 0.921 C. 0.003 D. 0.083

OpenStudy (anonymous):

This problem kind of sucks. Here we go: First, note that in probability problems, in general, x means "and," while + means "or." Now, on any given roll, the probability of rolling a 2 is 1/6. (duh--we're using a fair die with 6 numbers). So the probability of rolling anything else is 1-1/6=5/6. So, we are rolling a die 20 times. We need to get a 2 exactly 5 times, which means we need to get something other than a 2 exactly 15 times. Then, of exactly 20 rolls, we need 5 to be 2. However, note that these can be any 5 rolls, not necessarily the first 5 or the last 5 or any other specific order. So order doesn't matter and we use a combination, and since we're taking 5 rolls out of 20, that would be 20C5. So putting to gether these three parts yields: 20C5 x (1/6)^5 x (5/6)^15 Multiplying that out yields the right answer. http://answers.yahoo.com/question/index?qid=20070729220226AAUam3V

OpenStudy (anonymous):

A die is rolled 20 times and the number of twos that comes up is tallied. Find the probability of getting the given result. Exactly five twos. ----------------------- P(x=5) = 20C5(1/6)^2(5/6)^18 = 0.1294... http://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.165415.html

OpenStudy (anonymous):

thanks yahoo

OpenStudy (anonymous):

welcome))..........

OpenStudy (anonymous):

u understood